I have two brothers. I am eight years older than my younger brother. The difference
of my age from my younger brother’s age is twice the difference of my age from my big
brother’s age. We are now getting old, so the sum of our ages is now 164. How old am I?

Answer by ankor@dixie-net.com(22740)

Let a = your age
let b = big brother's age
let c = younger brother's age

I have two brothers. I am eight years older than my younger brother.
a = c + 8
or
c = a - 8
:
The difference of my age from my younger brother’s age is twice the difference of my age from my big brother’s age.
8 = 2(b - a)
8 = 2b - 2a
2a + 8 = 2b
divide by 2
a + 4 = b
:
We are now getting old, so the sum of our ages is now 164. How old am I?
a + b + c = 164
replace b with (a+4); replace c with (a-8)
a + (a+4) + (a-8) = 164
3a - 4 = 164
3a = 164 + 4
3a = 168
a = 168/3
a = 56 is his age
:
to check this
56 - 8 = 48 young brothers age
and
56 + 4 = 60 is old brother's age
they add up to:
56 + 60 + 48 = 164